OpenStudy (anonymous):

The variables x and y vary inversely. If x = 5/2 when y = 50, find an equation relating x and y. Help?

hartnn (hartnn):

inverse variation ----> xy=k k is constant of variation. when x=5/2, y=50, can u find the value of k ??

OpenStudy (anonymous):

So 5/2*50 = k. So 125 = k. and y = 125/x?

hartnn (hartnn):

absolutely correct! good :)

OpenStudy (anonymous):

i need help if x= 5/2 when y=50 find an equation relating x and y

OpenStudy (anonymous):

help me